OpenStudy (anonymous):

Can Any one help me with this limit ? \[\lim_{x \rightarrow 2}\frac{1}{x}\left( \frac{1}{x-2} -\frac{1}{2}\right)\]

OpenStudy (unklerhaukus):

\[\frac{1}{x-2}=\frac{1}{x-2}\times1\]\[=\frac{1}{x-2}\times\frac{x-2}{x-2}\]\[=\frac{x-2}{x^2-4x+4}\] hmm that didn't help

OpenStudy (anonymous):

... I think like your way .. its not work.

OpenStudy (unklerhaukus):

\[=\frac{1}{x-2}\times\frac{x+2}{x+2}\]\[=\frac{x+2}{x^2-4}\] damn didn't help either

OpenStudy (anonymous):

hahaha ... Its true ... I spend 30 mint by verifying these way .. you wrote it . its give true answer .

OpenStudy (anonymous):

its give true answer in solution .. but not this . way .. any way .. i don't want to lead ppl to wrong answer .. i just look4 it ^_^

OpenStudy (unklerhaukus):

\[\lim_{x\rightarrow2}\quad\frac{x-2}{x^2-4x+4}\rightarrow\frac 00\], so you can use l'Hôpital's rule

OpenStudy (anonymous):

so ? does this lead for certain answer ? if i took Hopital rule does that make answer 1/4 ?

OpenStudy (anonymous):

@UnkleRhaukus , thnx for trying :)

OpenStudy (foolaroundmath):

The limit does not exist. 1/(x-2) becomes inf when approached from right and becomes -inf when approached from left

OpenStudy (anonymous):

@FoolAroundMath , the Limit exist , and its answer 1/4 ,, i just need to verify it :)

OpenStudy (foolaroundmath):

http://www.wolframalpha.com/input/?i=lim%28x-%3E2%29+%281%2F%28x-2%29-1%2F2%29%281%2Fx%29 limit does not exist

OpenStudy (anonymous):

I want , or wish to believe it not exist ,, for that star for you , thnx

OpenStudy (unklerhaukus):

the graph is quite convincing